Find the sample variance of the data in the following frequency distribution.

X Frequency
10 2
15 2
20 4
25 2
total (=n) 10

A. 18.00

B. 25.00

C. 26.00

D. 28.89
Answer #1
X Frequency (f) f*x f*x^2
10 2 20 200
15 2 30 450
20 4 80 1600
25 2 50 1250
Total 10 180 3500

mean = \bar x = \frac{\sum fx}{\sum f} = \frac{180}{10} = 18

Sample varinace

mean = \bar x = \frac{\sum f_i(x_i -\bar x)^2}{n-1} \\ = \frac{2\times (10-18)^2+2\times (10-18)^2+2\times (10-18)^2+2\times (10-18)^2}{9} = 28.89

Answer :D. 28.89

​Please give thumbs up if you like the answer.